Software Engineering Python Problem! Please send in EXACT code to solve
Answer #1

Psudeo Code :-

Begin

1.Enter a number

2. repete step 3, 4 ,5 till number !=0

3. remender=number%10

4. append remender to the list

5. number=number/10

6. reverse the list

7. Print the list

End

Executable Code in Pyhon (Use python3 to execute):-

z = input()
a=int(z)
b = []

while(a!=0):
rem=int(a%10)
b.append (rem)
a=int(a/10)
b.reverse()
print (b)

Output :-

Input 1234 Output

Explanation-

lets take example number =456

rem =456%10 = 6

list=[6]

number=456/10 =45 (typecast to integer)

rem=45%10 = 5

list[6,5]

number =45/10 =4

rem=4%10 =4

list[6,5,4]

num=4/10 =0 (break loop here)

reversed list is [4,5,6]

that is desired output.